OpenStudy (anonymous):

a 2-kg metal ball moving at a speed of 3 m/s strikes a 1-kg wooden ball that is at rest. after the collision, the speed of the metal ball is 1 m/s. Assuming momentum is conserved, what is the speed of the wooden ball

OpenStudy (anonymous):

The momentum equation is p=m*v Where m=mass and v=velocity respectively. since momentum is conserved you can assume p initial=p final. so to find p initial you need to get the momentum of both balls before the collision. m*v(metal ball)+m*v(wooden ball)=p 2*3+1*0=p 6=p so then p final should also = 6 set the same equation = 6 but use your given speed of after the collision m*v(metal ball)+m*v(wooden ball)=p 2*1+1*X=6 2+X=6 X=4 So then the speed would be 4 m/s Hope this helps